A volume of 0.36 m3 of air initially at a temperature of 20oC and a 03 pressure of 4bar is compressed reversibly and isothermally to a final

volume of 0.06m3. Calculate the mass, the final pressure, work done and

heat transfer. The specific heat and gas constant can be taken as 1kJ/kgK

and 287 J/kg


1
Expert's answer
2022-01-28T09:28:01-0500

"m =\frac{P\times V\times M} {R\times T} = \frac{4\times 10^5 \times 0.36 \times 29} {287\times 293} =49.66" Kg

For Isothermal Process

"P_1V_1=P_2V_2"

"P2 = P_1\frac{V_1}{V_2} = 4\times \frac{0.36}{0.06} = 24 Bar"

work done

"W_{1-2} = p_1V_1" In "\frac{V_2}{V_1}"

"W_{1-2} =4\times10^5\times0.36" In "\frac{0.06}{0.36}=" 258.01Kw